# Active Generation for Image Classification

### v1

@@ Abstract
Recently, the growing capabilities of deep generative models have underscored their potential in enhancing image classification accuracy. However, existing methods often demand the generation of a disproportionately large number of images compared to the original dataset, while having only marginal improvements in accuracy. This computationally expensive and time-consuming process hampers the practicality of such approaches. In this paper, we propose to address the efficiency of image generation by focusing on the specific needs and characteristics of the model. With a central tenet of active learning, our method, named ActGen, takes a training-aware approach to image generation. It aims to create images akin to the challenging or misclassified samples encountered by the current model and incorporates these generated images into the training set to augment model performance. ActGen introduces an attentive image guidance technique, using real images as guides during the denoising process of a diffusion model. The model’s attention on class prompt is leveraged to ensure the preservation of similar foreground object while diversifying the background. Furthermore, we introduce a gradient-based generation guidance method, which employs two losses to generate more challenging samples and prevent the generated images from being too similar to previously generated ones. Experimental results on the CIFAR and ImageNet datasets demonstrate that our method achieves better performance with a significantly reduced number of generated images.
@@ 3.2 Text-conditioned guidance
In this paper, we directly use the public text-to-image diffusion models [35, 28] trained on large-scale text-image pairs as our generative models, with our innovative guidance methods introduced in inference process.
@@ 4.2 Attentive image guidance
$\tilde{\bm{x}}_{t-1}=\bm{x}_{t-1}+\bm{m}_{t}\odot\gamma_{t}(\bm{x}_{t-1}-\bm{x}_{t-1}^{(g)}),$ (8)
@@ 4.3 Gradient-based guidance
Adversarial samples. To exert further control over the difficulty of generation, we introduce a negative classification loss aimed at increasing the classification difficulty of synthetic images in accordance with the current model. Our approach involves decoding the image latent $\bm{x}_{t}$ in the diffusion model to obtain the original image $\bm{o}_{t}$. Subsequently, this image is fed into our current classification model $\Omega$ to obtain predicted logits $\Omega(\bm{o}_{t})$, which are then used to compute the negative cross-entropy (CE) loss with the corresponding label $y$, i.e.,
@@
Table 1: Top-1 accuracy on ImageNet classfication. SD denotes randomly generating samples with fixed prompt in our method. ***: Azizi et al. [1] uses Imagen [36], a generative model that is not open-source and is more powerful than the Stable Diffusion used in other methods. We reimplement Azizi et al. [1] in Table 5 for fairer comparisons.
@@ 5.4 Ablation study
Compare with Azizi et al. [1] on Stable Diffusion. One of our prior work, Azizi et al. [1], which generates images on ImageNet classes to improve the classification performance, leveraged the powerful Imagen [36] generative model. However, the model is not publicly available. For a fairer comparison, we implement Azizi et al. [1] on ImageNet with Stable Diffusion and generate the same number of images as our ActGen. As the results reported in Table 5, [1] on Stable Diffusion performs worse than the origin on Imagen, while our ActGen outperforms all of them.
Model Real only Azizi et al. [1] Azizi et al. [1] (SD) ActGen (ours)
ResNet-50 76.39 78.17 76.83 78.65
ViT-S/16 79.89 81.00 80.41 81.12
Training cost analysis. In ActGen, the generation of hard images and training of them introduce additional computational cost. Taking the training of ResNet-50 on ImageNet as an example. the traditional training with real images takes 9.6 GPU days on 32 NVIDIA V100 GPUs. While for ActGen, it takes 4.5 GPU days for generating images, resulting in 15.2 GPU days of total training time. However, compared to generating 10×10$\times$10 × of samples in previous method, which would cost $\sim$40 GPU days for generation and 2×2$\times$2 × time for training, our additional training cost is acceptable.
@@ 0.B.1 Few-shot classification on Pets dataset
0.B.2 More ablation studies

### v2

@@ Abstract
Recently, the growing capabilities of deep generative models have underscored their potential in enhancing image classification accuracy. However, existing methods often demand the generation of a disproportionately large number of images compared to the original dataset, while having only marginal improvements in accuracy. This computationally expensive and time-consuming process hampers the practicality of such approaches. In this paper, we propose to address the efficiency of image generation by focusing on the specific needs and characteristics of the model. With a central tenet of active learning, our method, named ActGen, takes a training-aware approach to image generation. It aims to create images akin to the challenging or misclassified samples encountered by the current model and incorporates these generated images into the training set to augment model performance. ActGen introduces an attentive image guidance technique, using real images as guides during the denoising process of a diffusion model. The model’s attention on class prompt is leveraged to ensure the preservation of similar foreground object while diversifying the background. Furthermore, we introduce a gradient-based generation guidance method, which employs two losses to generate more challenging samples and prevent the generated images from being too similar to previously generated ones. Experimental results on the CIFAR and ImageNet datasets demonstrate that our method achieves better performance with a significantly reduced number of generated images. Code is available at https://github.com/hunto/ActGen.
@@ 3.2 Text-Conditioned Guidance
In this paper, we directly use the public text-to-image diffusion models [35, 28] trained on large-scale text-image pairs as our generative models. Based on the fixed text condition a photo of <class>, we further propose additional guidance methods to control the generation in inference process.
@@ 4.2 Attentive Image Guidance
$\tilde{\bm{x}}_{t-1}=\bm{x}_{t-1}+\bm{m}_{t}\odot\gamma_{t}(\bm{x}_{t-1}^{(g)}-\bm{x}_{t-1}),$ (8)
@@ 4.3 Gradient-based Guidance
Adversarial samples. To exert further control over the difficulty of generation, we introduce a negative classification loss aimed at increasing the classification difficulty of synthetic images in accordance with the current model. Our approach involves denoising $\bm{x}_{t}$ to cleaned latent $\bm{x}_{0}$ with the diffusion model, then obtaining the original image $\bm{o}_{t}$. Subsequently, this image is fed into our current classification model $\Omega$ to obtain predicted logits $\Omega(\bm{o}_{t})$, which are then used to compute the negative cross-entropy (CE) loss with the corresponding label $y$, i.e.,
@@
Table 1: Top-1 accuracy on ImageNet classfication. SD denotes randomly generating samples with fixed prompt in our method. *: Azizi et al. [1] uses Imagen [36], a generative model that is not open-source and is more powerful than the Stable Diffusion used in other methods. We reimplement Azizi et al. [1] in supplementary material for fairer comparisons.
@@ 5.4 Ablation Study
Training cost analysis. In ActGen, the generation of hard images and training of them introduce additional computational cost. Taking the training of ResNet-50 on ImageNet as an example. the traditional training with real images takes 9.6 GPU days on 32 NVIDIA V100 GPUs. While for ActGen, it takes 4.5 GPU days for generating images, resulting in 15.2 GPU days of total training time. However, compared to generating 10×10$\times$10 × of samples in previous method, which would cost $\sim$40 GPU days for generation and 2×2$\times$2 × time for training, our additional training cost is acceptable. Meanwhile, comparing with Stable Diffusion v1.5, ActGen increases its VRAM of from 8.4G to 8.7G. Additionally, the generation times are comparable, with SD at 2.8 s/img and ActGen at 3.0 s/img.
@@ 6 Conclusion
Acknowledgements
This work was supported in part by the Australian Research Council under Projects DP210101859 and FT230100549.
@@ 0.B.1 Few-shot Classification on Pets Dataset
0.B.2 Compare with Azizi et al. [1] on Stable Diffusion
One of our prior work, Azizi et al. [1], which generates images on ImageNet classes to improve the classification performance, leveraged the powerful Imagen [36] generative model. However, the model is not publicly available. For a fairer comparison, we implement Azizi et al. [1] on ImageNet with Stable Diffusion and generate the same number of images as our ActGen. As the results reported in Table 9, [1] on Stable Diffusion performs worse than the origin on Imagen, while our ActGen outperforms all of them.
Model Real only Azizi et al. [1] Azizi et al. [1] (SD) ActGen (ours)
ResNet-50 76.39 78.17 76.83 78.65
ViT-S/16 79.89 81.00 80.41 81.12
We also compared our method with traditional training, original SD, and Azizi et al. [1] (excluding its cost of training generative model). As shown in Table 10, Azizi et al. [1] incurs over 3.8× training cost to traditional training, while our method increases traditional training by only 30% yet still outperforms Azizi et al. [1].
Table 10: Total generation and training time on ImageNet dataset.
Method Real only Azizi [1] SD Random Ours
Time (GPU hours) 384 $>1467$ 489 496
0.B.3 More Ablation Studies